OpenStudy (anonymous):

I have a question regarding complex numbers: u=2+i u*=2-i By considering the argument of (u/u*), or otherwise, prove that tan^-1(4/3)=2tan^-1(1/2)

OpenStudy (anonymous):

okay let's solve it what is value of (u/u*)? can you do it?

OpenStudy (anonymous):

Sorry for the late reply. Yes, I know the answer is (4+4i-1)/5 , I can't solve the rest.

OpenStudy (anonymous):

let's do it... your answer is correct... u/u*=\[(4+4i-1)/5 =(3+4i)/5\] \[=3/5+4/5i\] now what is argument of above complex no.? by comparing to standard form of complex no. z=x+yi we get x=3/5 and y=4/5 argument theta of (u/u*) is \[\theta=\tan^{-1} (y/x)\] \[\theta=\tan^{-1} (4/3)\] .................(1) now you can covert given complex nos.u and u* in polar form of (r,theta) [r=modulus of complex no. and theta is argument of complex no.] you get, \[u=(\sqrt{3},\theta _{1})\] and \[u*=(\sqrt{3},\theta _{2})\] where \[\theta _{1}=\tan ^{-1}(1/2)\] also \[\theta_{2}=\tan^{-1} (-1/2)=-\tan^{-1}(1/2) \] (as we know tan^-1(-x)=-tan^-1(x) ) now what we have to do.. we have to find (u/u*) now let's to this division in polar form now here is key when we do division in polar form the resultant complex no. has magnitude =magnitude of complex no . in numerator / magnitude of complex no. in denominator and resultant argument is=argument of complex no. in numerator - argument of complex no. in denominator that is if \[a=(r _{1},\theta _{1}) \] and \[b=(r _{2},\theta _{2})\] magnitude of complex no. (a/b)=(r1/r2) and argument is = (theta1 -theta2) now here we have a=u and b=u* and so magnitude of complex no (u/u*) is \[\sqrt{3}/\sqrt{3}=1\] and resultant argument of (u/u*) is \[\theta _{1}-\theta _{2}=\tan^{-1} (1/2)- [-\tan^{-1} (1/2)=\tan^{-1} (1/2)+\tan^{-1}(1/2)=2\tan^{-1} (1/2)\]..........(2) so from results (1) and (2) it is proved that tan^-1(4/3)=2tan^-1(1/2) !!! i hope this helps :)

OpenStudy (anonymous):

Thanks! The answer helped a lot!

OpenStudy (anonymous):

you are most welcome...! :)